# Spatial structure facilitates evolutionary rescue by cost-free drug resistance

### v1

@@ Abstract
Bacterial populations often have complex spatial structures, which can impact their evolution. Here, we study how spatial structure affects the evolution of antibiotic resistance in a bacterial population. We consider a minimal model of spatially structured populations where all demes (i.e., subpopulations) are identical and connected to each other by identical migration rates. We show that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, starting from a sensitive inoculum. Indeed, the bacterial population can be rescued if antibiotic resistant mutants appear and are present when drug is added, and spatial structure can impact the fate of these mutants and the probability that they are present. Specifically, if the mutation that provides resistance is neutral or effectively neutral, its probability of fixation is increased in smaller populations. This promotes local fixation of resistant mutants in the structured population, which facilitates evolutionary rescue by cost-free drug resistance. Once the population is rescued by resistance, migrations allow resistant mutants to spread in all demes. Our main results extend to the case where there are resistant mutants in the inoculum, and to more complex spatial structures. They also extend to resistant mutants that carry a fitness cost, although the timescales involved are longer.
@@
Antibiotic resistance is a major challenge, since bacteria tend to adapt to the drugs they are subjected to. Understanding what conditions facilitate or hinder the appearance and spread of resistance in a bacterial population is thus of strong interest. Most natural microbial populations have complex spatial structures. This includes host-associated microbiota, such as the gut microbiota. Here, we show that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, by promoting the presence of resistant bacteria. Indeed, neutral mutants giving resistance can take over small populations more easily than large ones, thanks to the increased importance of fluctuations in small populations. Resistant mutants can then spread to the whole structured population. Thus, population spatial structure can be a source of antibiotic treatment failure. This effect of spatial structure is generic and does not require environment heterogeneity.
@@ Introduction
Bacterial populations often have complex spatial structures. For pathogenic bacteria, each patient constitutes a different environment, connected through transmission events. Within a patient, different organs and tissues represent spatial environments between which bacteria can migrate [she_defining_2024]. At a smaller scale, bacteria in the gut microbiota, which can be a reservoir of drug resistance, live in different spatial environments, namely the digesta, the inter-fold regions and the mucus, and some of them form biofilms [donaldson_gut_2016, mccallum_gut_2023, chikina_attheright_2021], which are dense structures of cells enclosed in a polymer-based matrix [costerton_microbial_1995, watnick__biofilm_2000]. Thus, it is important to understand how the spatial structure of microbial populations impacts the evolution and spread of resistance. This question has been explored in epidemiological models in the case of viruses [Debarre07, Althouse13] and of bacteria [krieger_population_2020], and was recently addressed in controlled experiments, for specific antibiotic resistant bacteria [verdon_spatial_2024], but is still relatively under-explored. Spatial structure can lead to small effective population sizes, and small population sizes were recently experimentally shown to substantially impact the evolution and spread of resistance [Coates18, alexander_stochastic_2020]. More generally, the impact of spatial structure on the fate of mutants was studied theoretically [Wright31, Kimura64, maruyama70, maruyama74, Nagylaki80, Barton93, Whitlock97, Nordborg02, whitlock2003, Sjodin05, lieberman_evolutionary_2005, Houchmandzadeh11, allen_evolutionary_2017, marrec_toward_2021, yagoobi_fixation_2021, tkadlec_evolutionary_2023, abbara_frequent_2023] and experimentally [Kryazhimskiy12, Nahum15, france_relationship_2019, chakraborty_experimental_2023, kreger_role_2023], but generally in constant environments.
@@
Here, we ask how spatial structure impacts the fate of a bacterial population suddenly subjected to antibiotic treatment. Is the population eradicated by the drug, or does it survive by developing resistance? We address this question in a minimal model of population structure, composed of identical demes (i.e. subpopulations) connected to each other by identical migrations. This simple structure is known as the island model [Wright31, Kimura64], or the clique or fully-connected graph [lieberman_evolutionary_2005, marrec_toward_2021, abbara_frequent_2023]. We assume that the population initially only comprises drug-sensitive bacteria. Indeed, the initial diversity of a population of pathogenic bacteria in a host is often low, as many infections begin with a small infectious dose [gutierrez_virus_2012, lim_independent_2014, lythgoe_sarscov2_2021, woodward_gastric_2022, hoces_fitness_2023]. For simplicity, we assume that the environment is homogeneous, but note that environmental heterogeneities have a substantial effect on antibiotic resistance  [zhang_acceleration_2011, hermsen_onthe_2012, greulich_mutational_2012], and on evolutionary rescue in structured populations [Uecker14, Tomasini20, Tomasini22]. Specifically, we consider a perfect biostatic drug, which stops the growth of sensitive bacteria, but we also discuss the generalization of our results to other drug modes of action. When a well-mixed population is subjected to a treatment by a perfect biostatic drug (for a long enough time), the population gets extinct, except if resistant mutants are present when the drug is added [marrec_quantifying_2018, marrec_resist_2020, marrec_adapt_2020]. The population can thus be rescued by resistance. We study the impact of population spatial structure on this process. We also analyze how the time when drug is added impacts the fate of a spatially structured population.
@@
We show that spatial structure increases the probability that a bacterial population survives treatment by developing resistance through a neutral mutation. Specifically, for a given time of addition of the antibiotic, the survival probability of the population increases when the migration rate between demes is decreased, and when the number of demes is increased while keeping the same total population size. We show that this is due to the local fixation of resistant mutants in one or several demes. We study the composition of the population versus time before the addition of drug, and find that it is strongly impacted by spatial structure. We further study the time needed for resistant bacteria to colonize all demes once they have fixed in at least one of the demes. Finally, we show that our main results extend to the case where resistant mutants carry a cost, although the timescales are longer. They also extend to the case where resistant mutants are present in the inoculum, and to more complex spatial structures.
@@
We aim to assess the impact of spatial structure on the establishment of antibiotic resistance, in a minimal model where spatial structure is as simple as possible. Thus, we consider a spatially structured bacterial population comprising $D$ demes (i.e. subpopulations) on the nodes of a clique (i.e. a fully connected graph). This corresponds to the island population model [Wright31]. Each deme has the same carrying capacity $K$ (Figure 1, center). For comparison, we also consider a well-mixed population with the same total carrying capacity $KD$ (Figure 1, left), and a fully subdivided population composed of $D$ demes with carrying capacity $K$ without migrations between them (Figure 1, right). We model migrations between demes through a per capita migration rate $\gamma$, which is the same between all pairs of demes.
@@
Bacterial growth and resistance.
We consider two types of bacteria: sensitive (S) and resistant (R) ones. In the absence of antibiotics, the two types of bacteria have fitnesses $f_{S}$ and $f_{R}$, respectively, which set their maximal growth rates (in the exponential phase). They have the same death rate $g$. Thus, we implement selection on birth. However, it is straightforward to generalize our model to selection on death. We assume that growth is logistic in each deme. The growth rate of S (resp. R) individuals is $f_{S}\left(1-N/K\right)$ (resp. $f_{R}\left(1-N/K\right)$), where $N$ is the current population size of the deme considered. When S bacteria divide, their offspring can mutate to R with probability $\mu$. Here, we focus on the rare mutation regime, where $N\mu\ll 1$.
We mainly focus on neutral mutations, but we discuss the impact of a cost or of a benefit of resistance below and in SI Section 3.1. Considering neutral mutations is relevant because resistance mutations are often effectively neutral, e.g. at low drug concentrations. We thus set $f_{S}=f_{R}=1$, implying that the time unit in our system is set by the maximum growth rate of bacteria. We do not model further mutations or backmutations – note that the rate of backmutations was estimated to be one order of magnitude smaller than the rate of compensatory mutations [levin_compensatory_2000].
@@
As we aim to model the appearance of resistance, we start with a population comprising only S bacteria. In practice, we initialize each deme (resp. the well-mixed population) with a number of S bacteria representing 10% of $K$ (resp. of $KD$). Then, each deme (resp. the well-mixed population) quickly grows and reaches a steady-state size $K(1-g/f_{S})$ (resp. $KD(1-g/f_{S})$) around which it fluctuates. We choose this initial condition because it is appropriate to model the start of an infection [abel_analysis_2015]. Note that since $N\mu\ll 1$, it is very unlikely that mutations happen during this quick initial growth phase, whose timescale is set by $f_{S}=1$ [marrec_resist_2020]. Our results are thus robust to the initialization, as long as it is performed only with S individuals. We work in the regime where $g\ll f_{S}=1$, so that extinctions of demes are very unlikely in the absence of drug, and demes fluctuate around their steady-state size.
@@
We obtain analytical results from the Moran model [moran_random_1958] and from probability theory. Note that the Moran model assumes that population size is strictly fixed, which is not the case here. However, because we are in the regime where fluctuations around the steady-state size are small in the absence of drug, the Moran process then provides good approximations [marrec_resist_2020].
We perform stochastic simulations of the population evolution, using the Gillespie algorithm [gillespie_general_1976, gillespie_exact_1977]. The details of the simulation framework are described in the SI Section 2.1.
@@ Results
Figure 2A shows the survival probability versus the addition time of the antimicrobial for a spatially structured population with different values of the migration rate. We observe that survival probability increases with drug addition time. Indeed, R mutants are more likely to appear and fix in a population initially composed only by S bacteria if there is more time before drug addition. Moreover, we find that survival probability is higher when migration rate $\gamma$ is smaller.
Figure 2: Survival probability of a structured bacterial population subjected to biostatic antimicrobial. The survival probability is shown versus the drug addition time. Panel A: Spatially structured populations composed of $D=10$ demes with $K=100$ are considered, with different per capita migration rates $\gamma$. Results for the well-mixed population with same total size (“WMP”) are shown for reference. Panel B: Three levels of subdivision of a population with total carrying capacity $KD=1000$, into 5, 10, and 20 demes, are shown for two values of $\gamma$. Results for the well-mixed population (“WMP”) are shown for reference. In all cases, we determined the survival probability from the fraction of $10^{4}$ simulations where the population survived after drug application. The vertical dash-dotted line denotes the average time of appearance of a locally successful mutant in the fastest deme (Equation (2)). The vertical dotted line represents the average appearance time of a successful mutant in a well-mixed population (Equation (1)). Thin lines connecting data points are guides for the eye. The horizontal gray solid line in both panels is the probability of presence of R mutants belonging to a lineage destined to go extinct, calculated from Equation (S16). Parameter values: $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=1$, $g=0.1$, $\mu=10^{-5}$.
Figure 2B shows the impact of varying the degree of subdivision of a population of fixed total size, but composed of a different number of demes. We observe that higher subdivision yields higher probabilities of survival.
@@
To rationalize our results on the impact of spatial structure on the ability of a microbial population to survive the addition of biostatic drug, we focus on rescue by resistance. R mutants can arise in our population with rate $N_{tot}\mu g$, where $N_{tot}$ is the total number of individuals in the population, $\mu$ the mutation probability upon division and $g$ the death rate, which equals the division rate at steady state. All the populations we considered in Figure 2 have the same total steady-state size $N_{tot}=KD(1-g/f_{S})$ in the absence of drug. Thus, the mutant supply is not impacted by spatial structure. However, the fate of mutants can be impacted by spatial structure. Most R mutants that appear give rise to lineages that go extinct, whatever the structure. The survival time of their lineages is not impacted by population size. Let us now consider mutants that fix. Since we consider neutral mutants, their fixation probability is $1/N$ in a population of size $N$. In spatially structured populations, fixation can occur locally within a deme, and then it is governed by the smaller steady-state population size $N=K(1-g/f_{S})$ of a deme. If migrations are rare enough, once R individuals have fixed in a deme, they survive there, and can rescue the population.
@@
To understand the impact of local mutant fixation on population survival, let us focus on successful mutants, i.e. those that give rise to a lineage that fixes. The average time of appearance $\langle t_{af\,W}\rangle$ of a successful mutant in a well-mixed population of size $N$ is given by:
$\langle t_{af\,W}\rangle=t_{app}\times\frac{1}{p_{fix}}=\frac{1}{N\mu g}\times N=\frac{1}{\mu g}\,,$ (1)
where $t_{app}=1/(N\mu g)$ is the average appearance time of a mutant and $p_{fix}=1/N$ is its fixation probability. Recall that we are considering neutral mutants (see above, and SI Section 1.1). Importantly, this result does not depend on population size $N$. Therefore, it holds both for the well-mixed population and for each deme in a structured population.
@@
In a structured population, the first mutant that fixes locally in one deme may rescue the population by its resistance. The appearance of such a locally successful mutant is a Poisson process with rate $\mu g$ (see Equation (1)). Thus, the average appearance time of a locally successful mutant in the fastest deme is (see SI Section 1 for details):
$\langle t_{af\,F}\rangle=\frac{1}{D\mu g}\,.$ (2)
Crucially, this is $D$ times faster than the appearance of a successful mutant in a well-mixed population (which is given by Equation (1)).
@@
The addition of a biostatic antimicrobial (leading to $f_{S}=0$, $f_{R}=1$ when drug is added) should have a different effect depending how the addition time $t_{add}$ compares to the two timescales defined in Equations (1) and (2). Specifically, we predict that:
@@
  • If $t_{add}\ll\langle t_{af\,F}\rangle$, the bacterial population is likely to be eradicated, whatever its structure.
@@
  • If $\langle t_{af\,F}\rangle\ll t_{add}\ll\langle t_{af\,W}\rangle$, it is likely that a locally successful mutant has appeared. This rescues a structured population. However, it is likely that no successful mutant has appeared in the well-mixed population yet. A well-mixed population is thus expected to go extinct.
@@
  • If $t_{add}\gg\langle t_{af\,W}\rangle$, both the well-mixed and the structured population are expected to survive.
@@
In Figure 2, we show the two timescales defined in Equations (1) and (2). We observe that the most substantial impact of structure on survival probabilities is observed when $\langle t_{af\,F}\rangle<t_{add}<\langle t_{af\,W}\rangle$. This is fully in line with our theoretical analysis based on timescale comparisons. Furthermore, we observe that all populations survive if $t_{add}\gg\langle t_{af\,W}\rangle$. Finally, if $t_{add}\ll\langle t_{af\,F}\rangle$, the population is eradicated in most cases. It can nevertheless be rescued by non-successful mutant lineages (see also Figure S3). Thus, the probability of treatment survival is given by the probability that non-successful mutants are present when the drug is added, see Equation (S16). Our simulation results are in good agreement with this analytical prediction, see Figure 2.

### v2

# Spatial structure facilitates evolutionary rescue by drug resistance

@@ Abstract
Bacterial populations often have complex spatial structures, which can impact their evolution. Here, we study how spatial structure affects the evolution of antibiotic resistance in a bacterial population. We consider a minimal model of spatially structured populations where all demes (i.e., subpopulations) are identical and connected to each other by identical migration rates. We show that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, starting from a sensitive inoculum. Specifically, the bacterial population can be rescued if antibiotic resistant mutants appear and are present when drug is added, and spatial structure can impact the fate of these mutants and the probability that they are present. Indeed, the probability of fixation of neutral or deleterious mutations providing drug resistance is increased in smaller populations. This promotes local fixation of resistant mutants in the structured population, which facilitates evolutionary rescue by drug resistance in the rare mutation regime. Once the population is rescued by resistance, migrations allow resistant mutants to spread in all demes. Our main result that spatial structure facilitates evolutionary rescue by antibiotic resistance extends to more complex spatial structures, and to the case where there are resistant mutants in the inoculum.
@@
Antibiotic resistance is a major challenge, since bacteria tend to adapt to the drugs they are subjected to. Understanding what conditions facilitate or hinder the appearance and spread of resistance in a bacterial population is thus of strong interest. Most natural microbial populations have complex spatial structures. This includes host-associated microbiota, such as the gut microbiota. Here, we show that spatial structure can facilitate the survival of a bacterial population to antibiotic treatment, by promoting the presence of resistant bacteria. Indeed, neutral or deleterious mutants giving resistance can take over small populations more easily than large ones, thanks to the increased importance of fluctuations in small populations. Resistant mutants can then spread to the whole structured population. Thus, population spatial structure can be a source of antibiotic treatment failure. This effect of spatial structure is generic and does not require environment heterogeneity.
@@ Introduction
Bacterial populations often have complex spatial structures. For pathogenic bacteria, each patient constitutes a different environment, connected through transmission events. Within a patient, different organs and tissues represent spatial environments between which bacteria can migrate [she_defining_2024]. At a smaller scale, bacteria in the gut microbiota, which can be a reservoir of drug resistance, live in different spatial environments, namely the digesta, the inter-fold regions and the mucus, and some of them form biofilms [donaldson_gut_2016, mccallum_gut_2023, chikina_attheright_2021], which are dense structures of cells enclosed in a polymer-based matrix [costerton_microbial_1995, watnick__biofilm_2000]. Thus, it is important to understand how the spatial structure of microbial populations impacts the evolution and spread of resistance. This question has been explored in epidemiological models in the case of viruses [Debarre07, Althouse13] and of bacteria [krieger_population_2020], and was recently addressed in controlled experiments, for specific antibiotic resistant bacteria [verdon_spatial_2024], but is still relatively under-explored. Spatial structure can lead to small effective population sizes, and small population sizes were recently experimentally shown to substantially impact the evolution and spread of resistance [Coates18, alexander_stochastic_2020]. More generally, the impact of spatial structure on the fate of mutants has been studied theoretically [Wright31, Kimura64, maruyama70, maruyama74, Nagylaki80, slatkin_fixation_1981, Barton93, Whitlock97, Nordborg02, whitlock2003, Sjodin05, lieberman_evolutionary_2005, Houchmandzadeh11, hauert_fixation_2014, allen_evolutionary_2017, marrec_toward_2021, yagoobi_fixation_2021, tkadlec_evolutionary_2023, abbara_frequent_2023] and experimentally [Kryazhimskiy12, Nahum15, france_relationship_2019, chakraborty_experimental_2023, kreger_role_2023], but generally in constant environments. Furthermore, theoretical investigations have mainly focused on the fate of one mutant lineage, specifically on mutant fixation probability and fixation time.
@@
Here, we ask how spatial structure impacts the overall fate of a bacterial population suddenly subjected to antibiotic treatment. Is the population eradicated by the drug, or does it survive by developing resistance? We address this question in a minimal model of population structure, composed of identical demes (i.e. subpopulations) connected to each other by identical migrations. This simple structure is known as the island model [Wright31, Kimura64], or the clique or fully-connected graph [lieberman_evolutionary_2005, marrec_toward_2021, abbara_frequent_2023]. We assume that the population initially only comprises drug-sensitive bacteria. Indeed, the initial diversity of a population of pathogenic bacteria in a host is often low, as many infections begin with a small infectious dose [gutierrez_virus_2012, lim_independent_2014, lythgoe_sarscov2_2021, woodward_gastric_2022, hoces_fitness_2023]. For simplicity, we assume that the environment is homogeneous, but note that environmental heterogeneities have a substantial effect on antibiotic resistance  [zhang_acceleration_2011, hermsen_onthe_2012, greulich_mutational_2012], and on evolutionary rescue in structured populations [Uecker14, Tomasini20, Tomasini22]. Specifically, we consider a perfect biostatic drug, which stops division of sensitive bacteria, but we also discuss the generalization of our results to other drug modes of action. When a well-mixed population is subjected to a treatment by a perfect biostatic drug (for a long enough time), the population gets extinct, except if resistant mutants are present when the drug is added [marrec_quantifying_2018, marrec_resist_2020, marrec_adapt_2020]. The population can thus be rescued by resistance. We study the impact of population spatial structure on this process. We also analyze how the time when drug is added impacts the fate of a spatially structured population.
@@
We show that spatial structure increases the probability that a bacterial population survives treatment by developing resistance through a neutral or deleterious mutation. Specifically, for a given time of addition of the antibiotic, the survival probability of the population increases when the migration rate between demes is decreased, and when the number of demes is increased while keeping the same total population size. We show that this is due to the local fixation of resistant mutants in one or several demes. We study the composition of the population versus time before the addition of drug, and find that it is strongly impacted by spatial structure. We discuss the parameter regimes where spatial structure facilitates rescue by resistance. We further study the time needed for resistant bacteria to colonize all demes after drug addition. Finally, we show that our main results regarding the impact of spatial structure on population survival to drug extend to more complex spatial structures, and to the case where resistant mutants are present in the inoculum.
@@
We aim to assess the impact of spatial structure on the establishment of antibiotic resistance, in a minimal model where spatial structure is as simple as possible. Thus, we consider a spatially structured bacterial population comprising $D$ demes (i.e. subpopulations) on the nodes of a clique (i.e. a fully connected graph). This corresponds to the island population model [Wright31]. Each deme has the same carrying capacity $K$ (Fig 1, center). For comparison, we also consider a well-mixed population with the same total carrying capacity $DK$ (Fig 1, left), and a fully subdivided population composed of $D$ demes with carrying capacity $K$ without migrations between them (Fig 1, right). We model migrations from one deme to another through a per capita migration rate $\gamma$, which is the same between all pairs of demes.
@@
Bacterial division, death and resistance.
We consider two types of bacteria: sensitive (S) and resistant (R) ones. In the absence of antibiotics, the two types of bacteria have fitnesses $f_{S}$ and $f_{R}$, respectively, which set their maximal division rates (in the exponential phase). They have the same death rate $g$. Thus, we implement selection on birth. However, it is straightforward to generalize our model to selection on death. We assume that growth is logistic in each deme. The division rate of S (resp. R) individuals is $f_{S}\left(1-N/K\right)$ (resp. $f_{R}\left(1-N/K\right)$), where $N$ is the current population size of the deme considered. When S bacteria divide, their offspring can mutate to R with probability $\mu$. Here, we focus on the rare mutation regime, where $DK\mu\ll 1$.
We set $f_{S}=1$, implying that the time unit in our system is set by the maximum division rate of sensitive bacteria. Because antibiotic resistance often carries a fitness cost in the absence of drug [Andersson99, Andersson10], we write $f_{S}=1-\delta$, where $\delta$ represents the cost of resistance and satisfies $0\leq\delta\ll 1$.
We do not model further mutations or backmutations – note that the rate of backmutations was estimated to be one order of magnitude smaller than the rate of compensatory mutations [levin_compensatory_2000].
@@
As we aim to model the appearance of resistance, we start with a population comprising only S bacteria. In practice, we initialize each deme (resp. the well-mixed population) with a number of S bacteria representing 10% of $K$ (resp. of $DK$). Then, each deme (resp. the well-mixed population) quickly grows and reaches a steady-state size $N^{*}=K(1-g/f_{S})$ (resp. $DN^{*}=DK(1-g/f_{S})$) around which it fluctuates. We choose this initial condition because it is appropriate to model the start of an infection [abel_analysis_2015]. Note that since $DK\mu\ll 1$, it is very unlikely that mutations happen during this quick initial population growth phase, whose timescale is set by $f_{S}=1$ [marrec_resist_2020]. Our results are thus robust to the initialization, as long as it is performed only with S individuals.
We work in the regime where $K\gg 1$ and $g\ll f_{S}=1$. Therefore, extinctions of demes are extremely slow in the absence of drug, and these events can be neglected here, as their timescales are much longer than all those we will discuss. (Specifically, the extinction of a deme with $K=100$ starting with 10 S bacteria, with $g=0.1$ and $f_{S}=1$, see Ref. [marrec_adapt_2020], would take an average time $2.5\times 10^{60}$ in the absence of drug.) Furthermore, in this regime, demes fluctuate weakly around their steady-state size in the absence of drug.
@@
We obtain analytical results from the Moran model [moran_random_1958] and from probability theory. Note that the Moran model assumes that population size is strictly fixed, which is not the case here. However, because we are in the regime where fluctuations around the steady-state size are small in the absence of drug, the Moran process provides good approximations (see Ref. [marrec_resist_2020]).
We perform stochastic simulations of the population evolution, using the Gillespie algorithm [gillespie_general_1976, gillespie_exact_1977]. The details of the simulation framework are described in the Supplementary Appendix Section 2.1.
@@ Results
Fig 2A shows the survival probability versus the time $t_{\textrm{add}}$ at which antimicrobial is added, for a spatially structured population with different values of the migration rate, in the case where resistant mutants carry no fitness cost, i.e. $\delta=0$. At the addition time $t_{\textrm{add}}$ of antimicrobial, the fitness of S bacteria switches from $f_{S}=1$ to $f_{S}=0$, while that of R bacteria remains the same. We observe that survival probability increases with $t_{\textrm{add}}$. Indeed, R mutants are more likely to appear and fix in the population, which is initially composed only of S bacteria, if there is more time before drug addition. Moreover, we find that survival probability is higher when migration rate $\gamma$ is smaller. In Supplementary Appendix Section 3.1, we further observe that subdividing a population with fixed total size into more numerous demes yields higher probabilities of survival.
Figure 2: Survival probability of a structured bacterial population subjected to biostatic antimicrobial. The survival probability is shown versus the drug addition time. Panel A: Spatially structured populations composed of $D=10$ demes with $K=100$ are considered, with different per capita migration rates $\gamma$. Results for the well-mixed population with same total size (“WMP”) are shown for reference. Markers report simulation results (each obtained from $10^{4}$ replicate simulations). The vertical dash-dotted line denotes the average time of appearance of a locally successful mutant in the fastest deme (Eq 2). The vertical dotted line represents the average appearance time of a successful mutant in a well-mixed population of size $DN^{*}$ (Eq 1). Solid lines for the fully subdivided ($\gamma=0$) and well-mixed populations are analytical predictions from Eq 4. Thin solid lines connecting markers are guides for the eye. Panel B: Same as in Panel A, except that mutants have a cost of resistance $\delta=10^{-2}$. Parameter values: $f_{S}=1$ without drug, $f_{S}=0$ with drug, $f_{R}=1$ (Panel A), $f_{R}=0.99$ (Panel B), $g=0.1$, $\mu=10^{-5}$.
Fig 2B shows the survival probability versus the drug addition time $t_{\textrm{add}}$ when resistant mutants have a small cost of resistance $\delta\ll 1$. We observe an even stronger impact of spatial structure on population survival than in the neutral case. Indeed, in the range of $t_{\textrm{add}}$ considered, the survival probability is roughly constant for the well-mixed population, while it strongly increases with $t_{\textrm{add}}$ for populations that are fully subdivided or have small migration rates. With the parameter choices of Fig 2B, resistant mutants are significantly deleterious in a well-mixed population ($DK\delta\gg 1$), while the impact of the cost is weaker within a deme ($K\delta=1$). Accordingly, the well-mixed population behaves in a very different way in Figs 2A and 2B, while the fully subdivided one behaves similarly.
Note that the parameter values of Fig 2 were chosen to be in the rare mutation regime ($DK\mu\ll 1$, see “Models and methods”), and for computational tractability. The parameter regimes where spatial structure facilitates rescue by resistance will be determined in a dedicated section below, and the connection to realistic values will be made in the Discussion.
@@
Spatial structure allows local mutant fixation.
@@
To rationalize our results on the impact of spatial structure on the ability of a microbial population to survive the addition of biostatic drug, we focus on rescue by resistance. Indeed, recall that we start from a sensitive inoculum. As mutations occur at division, R mutants arise with rate $N^{*}_{\textrm{tot}}\mu g$, where $N^{*}_{\textrm{tot}}$ is the steady-state total number of individuals in the population, $\mu$ the mutation probability upon division and $g$ the death rate, which equals the division rate at steady state. All the populations we considered in Fig 2 have the same total steady-state size $N^{*}_{\textrm{tot}}=DK(1-g/f_{S})$ in the absence of drug. Thus, the mutant supply is not impacted by spatial structure. However, the fate of mutants can be impacted by spatial structure.
Most R mutants that appear give rise to lineages that go extinct in the absence of drug, whatever the structure. Mutant lineages destined for extinction in the absence of drug can rescue the population if they are present when drug is added [marrec_resist_2020]. Indeed, once drug is added, S bacteria decay, and R bacteria can proliferate thanks to this reduced competition. However, this scenario remains unlikely in the rare mutation regime ($DK\mu\ll 1$), and is not strongly impacted by the value of $t_{\textrm{add}}$ or by population structure (see Supplementary Appendix Section 1.1). Thus, it cannot explain the impact of spatial structure on rescue by resistance that we observe in Fig 2.
Let us now consider mutants that fix. Crucially, in spatially structured populations, fixation can occur locally within a deme. If migrations are rare enough, once R individuals have fixed in a deme, they survive there, provided that migrations are rare enough that we can neglect a subsequent migration and fixation of S bacteria in these R demes. Upon the addition of drug, these locally successful mutants rescue the population. R bacteria can then spread by migration to the whole population.
Timescales of appearance of mutants that locally fix.
@@
To understand the impact of local mutant fixation on population survival, let us focus on successful mutants, i.e. those that give rise to a lineage that fixes. The average time $\langle t_{af\,W}(N^{*})\rangle$ of appearance (“$a$”) of a mutant that fixes (“$f$”) in a well-mixed population (“$W$”) of fixed size $N^{*}$ is given by [desai_beneficial_2007, weissman_rate_2009]:
$\langle t_{af\,W}(N^{*})\rangle=\langle t_{\textrm{app}}(N^{*})\rangle\times\frac{1}{p_{\textrm{fix}}(N^{*})}=\frac{1}{N^{*}\mu g\,p_{\textrm{fix}}(N^{*})}\,,$ (1)
where $\langle t_{\textrm{app}}(N^{*})\rangle=1/(N^{*}\mu g)$ is the average appearance time of a mutant, and $p_{\textrm{fix}}(N^{*})$ is its fixation probability, which can be expressed in the Moran model [ewens_mathematical_2004], see Supplementary Appendix Section 1.2. Note that here, we are interested in the time until a successful R mutant appears [desai_beneficial_2007, weissman_rate_2009], and not in its fixation time in a deme or in the full structured population [slatkin_fixation_1981, whitlock2003, hauert_fixation_2014].
@@
In a structured population, the first mutant that fixes locally in one deme may rescue the population by its resistance. Neglecting the small fluctuations of demes around their steady-state size $N^{*}=K(1-g/f_{S})$, the appearance of such a locally successful mutant can be approximated by a Poisson process with rate $1/\langle t_{af\,W}(N^{*})\rangle$ (see Eq 1), and appearance time is exponentially distributed. Thus, the average appearance time of a locally successful mutant in the fastest (“$F$”) deme among $D$ demes of steady-state size $N^{*}$ is (see Supplementary Appendix Section 1.4 for details):
$\langle t_{af\,F}(N^{*},D)\rangle=\frac{\langle t_{af\,W}(N^{*})\rangle}{D}=\frac{1}{DN^{*}\mu g\,p_{\textrm{fix}}(N^{*})}\,.$ (2)
Crucially, a locally successful R mutant appears in the fastest deme of a structured population faster on average than a successful R mutant in a well-mixed population with the same total size. Indeed, Eqs 1 and 2 give
$\frac{\langle t_{af\,F}(N^{*},D)\rangle}{\langle t_{af\,W}(DN^{*})\rangle}=\frac{p_{\textrm{fix}}(DN^{*})}{p_{\textrm{fix}}(N^{*})}<1\,,$ (3)
for all resistance costs $\delta\geq 0$. This ratio is smaller than 1 because neutral or deleterious mutations fix more easily in small populations than in large ones, due to the increased effect of genetic drift, i.e. of fluctuations due to finite-size effects. For neutral R mutants ($\delta=0$), we have $p_{\textrm{fix}}(N^{*})=1/N^{*}$, so the ratio in Eq 3 is equal to $1/D$. Thus, a locally successful mutant appears in a structured population $D$ times faster than a successful mutant in a well-mixed population with same total size. For R mutants with a nonzero cost of resistance $\delta>0$, the ratio in Eq 3 is even smaller, as mutant fixation is more strongly suppressed in larger population. See Supplementary Appendix Section 3.2 for details.
@@
Adding a biostatic antimicrobial (leading to $f_{S}=0$ and $f_{R}=1-\delta$ when drug is added) should thus have a different effect depending on the drug addition time $t_{\textrm{add}}$:
@@
  • If $t_{\textrm{add}}\ll\langle t_{af\,F}(N^{*},D)\rangle$, the bacterial population is likely to be eradicated, whatever its structure.
@@
  • If $\langle t_{af\,F}(N^{*},D)\rangle\ll t_{\textrm{add}}\ll\langle t_{af\,W}(DN^{*})\rangle$, it is likely that a locally successful mutant has appeared. This rescues a fully subdivided population (with no migrations). However, it is likely that no successful mutant has appeared in a well-mixed population yet. A well-mixed population is thus expected to go extinct.
@@
  • If $t_{\textrm{add}}\gg\langle t_{af\,W}(DN^{*})\rangle$, both a well-mixed and a fully subdivided population are expected to survive.
@@
In Fig 2A, where neutral mutants are considered, we show the two timescales compared in Eq 3. We observe that the most substantial impact of structure on survival probabilities is observed when $\langle t_{af\,F}(N^{*},D)\rangle<t_{\textrm{add}}<\langle t_{af\,W}(DN^{*})\rangle$. This is fully in line with our theoretical analysis based on timescale comparisons. Furthermore, we observe that all populations survive if $t_{\textrm{add}}\gg\langle t_{af\,W}(DN^{*})\rangle$. Finally, if $t_{\textrm{add}}\ll\langle t_{af\,F}(N^{*},D)\rangle$, the population is eradicated in most cases. It can nevertheless be rescued by mutant lineages that were destined for extinction in the absence of drug. Thus, the probability of treatment survival is then given by the probability that non-successful mutants are present when the drug is added, see Eq S4.
In Fig 2B, we consider mutants with moderate fitness cost $\delta=1/K$. The average appearance time $\langle t_{af\,F}(N^{*},D)\rangle$ of a successful mutant in the fastest deme is shown (cf. Eq 2 and Eq S22) and is of the same order of magnitude as in the neutral case of Fig 2A. By contrast, these mutants are significantly deleterious in the well-mixed population since $DK\delta=D\gg 1$. Hence, $\langle t_{af\,W}(DN^{*})\rangle$ is not shown in Fig 2B, as it evaluates to $2.2\times 10^{9}$, which is about $10^{4}$ times longer than $\langle t_{af\,F}(N^{*},D)\rangle$. The impact of spatial structure is even stronger in this case than for neutral mutants.
Analytical expression of the survival probability.
To go beyond our comparison of timescales, let us express the probability $p_{\textrm{surv}}(t_{\textrm{add}})$ that a population survives the addition of drug at time $t_{\textrm{add}}$, in a well-mixed population or in a fully subdivided population with no migrations. For this, we approximate deme size as constant, and we consider the two distinct mechanisms that can yield population survival, namely rescue by a (locally) successful mutant and rescue by the presence of the lineage of a mutant that was destined for extinction in the absence of drug. This leads to:
$p_{\textrm{surv}}(t_{\textrm{add}})=p_{\textrm{succ}}(t_{\textrm{add}})+\left[1-p_{\textrm{succ}}(t_{\textrm{add}})\right]p_{\textrm{pres}}\,,$ (4)
where $p_{\textrm{succ}}(t_{\textrm{add}})$ is the probability that a (locally) successful mutant has appeared by $t_{\textrm{add}}$, while $p_{\textrm{pres}}$ is the probability of presence of a lineage of R mutants destined to go extinct in the absence of drug. In the Supplementary Appendix, we calculate $p_{\textrm{pres}}$ in Section 1.1 (see Eq S4) and $p_{\textrm{succ}}(t_{\textrm{add}})$ in Section 1.6. The specific form of $p_{\textrm{succ}}(t_{\textrm{add}})$ is given by Eqs S28 and S29 for a well-mixed population, and by Eqs S31 and S32 for a fully subdivided population (with and without fitness cost, respectively).
Fig 2 shows that our simulation results are in very good agreement with the analytical predictions from Eq 4, both for the well-mixed population and for the fully subdivided population.